# Split Two-Periodic Aztec Diamond

### v4

@@ 4.4. Local Asymptotics of $I^{\alpha}_{2,1}$
From Lemma 4.2, we know that $\psi_{1}(z)$ has a maximum at $z_{1}^{*}$, so $\gamma_{0,1}$ should be deformed to the steepest ascent contour through this point $\gamma_{s,1}$. After this deformation, we only need to show that
@@ 5. An Intermediate Correlation Kernel via a Non-Intersecting Paths Process
In order to prove the intermediate correlation kernel presented in Lemma 3.1, we must first describe the relationship between the Aztec diamond and the non-intersecting paths process presented in [3]. This explanation can be found in Section 5.1. These results are not new, the bijection was first detailed in [3, Section 5] and then further described in [6]. Once we introduce this process, the formulation of the kernel follows from a slight modification of [3, Theorem 5.2]. This work is done in Section 5.2.
@@ 5.1. Non-Intersecting Paths Process
Recall the definition of the Aztec diamond graph given by equations (1)-(3). From this definition, we would like to label the four types of edges found in the Aztec diamond graph. We call an edge of the form $((2j+1,2k),(2j+2,2k+1))$ an east edge, an edge of the form $((2j+1,2k),(2j,2k+1))$ a west edge, an edge of the form $((2j+1,2k),(2j+2,2k-1))$ a south edge, and an edge of the form $((2j+1,2k),(2j,2k-1))$ a north edge. The different types of edges are depicted in Figure 16.
@@ 5.1.1. The Aztec Diamond to DR Paths
We will start by relating the Aztec diamond graph to the DR lattice paths model, which was first introduced in [15]. Under an appropriate choice of weighting this model is equivalent to the Aztec diamond. We will explain this bijection here. To start, we will describe the DR graph of size $n$. The DR graph of size $n$ has vertices given by,
@@
Figure 16. On the left is a covering of the Aztec diamond of size $n=4$. East, west, south, and north edges are depicted in blue, red, yellow, and green respectively. The corresponding paths on the DR graph are shown in the figure on the right.
@@ 5.1.2. The Tower Aztec Diamond Graph
The tower Aztec diamond graph was first described in [3] and later named in [6]. Informally, the $(n,p)$ tower Aztec diamond graph consists of two Aztec diamonds, one of size $n$ and the other of size $n-1$, stitched together by a strip of the rotated square grid of length $p$. An example of the tower Aztec diamond graph is show in Figure 17. The $(n,p)$ tower Aztec diamond graph has the vertex sets,
@@
(97) $E^{\text{Tow}}_{n,p}=\big{\{}((2j+1,2k),(2j+2,2k+1)):0\leq j\leq n-1,\,0\leq k\leq n-1\text{ or }\\ 0\leq j\leq n-2,\,-p-n+1\leq k\leq-1\big{\}}\\ \cup\big{\{}((2j+1,2k),(2j,2k+1)):0\leq j\leq n-1,\,-p\leq k\leq n-1\text{ or }\\ 0\leq j\leq n-2,\,-p-n+1\leq k\leq-1-p\big{\}}\\ \cup\big{\{}((2j+1,2k),(2j+2,2k-1)):0\leq j\leq n-1,\,1\leq k\leq n\text{ or }\\ 0\leq j\leq n-2,\,1-p-n\leq k\leq n\big{\}}\\ \cup\big{\{}((2j+1,2k),(2j,2k-1)):0\leq j\leq n-1,\,1-p\leq k\leq n\text{ or }\\ 1\leq j\leq n-1,1-p-n\leq k\leq n\big{\}}$
One should realize that it is not possible to have a dimer configuration on the tower Aztec diamond where one dimer contains both a vertex from the corridor and a vertex from either of the appending Aztec diamonds. In fact, there is only one possible dimer configuration on the corridor. Thus the statistics of the tower Aztec diamond can be simplified to the statistics on the two appending Aztec diamond. See [6] for more detail.
@@ 5.1.3. Tower Aztec Diamond to Paths Process
The tower DR path process is nearly the process described in [3]. To get exactly the process in their work, we should change the diagonal edges of the DR tower graph into a combination of a horizontal edge and diagonal edge, as depicted in Figure 18. The additional horizontal edges will have weight one and thus not effect the statistics of the model. Once again we consider $n+p$ non intersection paths which start on the left at the top $n+p$ vertices and end on the right at the bottom $n+p$ vertices.
Thus if we send the length of the corridor to infinity, the correlation kernel for the top part of the model,referred to as $K_{top}$, in [3]) is also a correlation kernel of the Aztec diamond of size $n$. This correlation kernel is stated explicitly in [3, Theorem 3.1].
@@
Introducing this parameter $a$ is necessary in order for us to directly apply [3, Theorem 3.1]. This is because the theorem does not permit singularities on the unit circle. We will call this weighting a split biased two-periodic weighting. Ultimately, we will send $a\to 1$ to obtain results for our model of interest. To compute the kernel, we will need to utilize the Weiner-Hopf factorization,
@@
For a full description of the derivation and properties of the matrices $\tilde{\phi}_{-,a,N}(z)$ and $\tilde{\phi}_{+,a,N}(z)$ see [3, Section 3]. It is important to note that $\tilde{\phi}_{+,a,N}(z)^{\pm 1}$ is analytic for $|z|<1$ and continuous for $|z|\leq 1$, while $\tilde{\phi}_{-,a,N}(z)^{\pm 1}$ is analytic for $|z|>1$ and continuous for $|z|\geq 1$. Moreover, the factorization commutes with the limit $a\to 1$.
@@ 5.2.1. Proof when $m^{\prime}\leq\nicefrac{{N}}{{2}}$
From [3, Theorem 3.1] we may immediately write the correlation kernel for the split biased two-periodic Aztec diamond.
@@
(104) $\phi_{m,a,N}(z)=\begin{cases}\Phi_{\alpha,a}(z)^{m}&\text{if }m\leq\nicefrac{{N}}{{2}}\\ \Phi_{\alpha,a}(z)^{\frac{N}{2}}\Phi_{\beta,a}(z)^{m-\frac{N}{2}}&\text{if }m>\nicefrac{{N}}{{2}}\end{cases}$
The contour $\gamma_{a}$ is a simple closed curve around $a^{2}$, while the contour $\gamma_{0,1,a}$ is a simple closed curve containing 00, $1$, and $\gamma_{a}$. To get the correlation kernel for the split two-periodic Aztec diamond, we wish to take the limit as $a\to 1$. The problem is the double contour integral in equation (103) is singular at both $a^{2}$ and $a^{-2}$ (with respect to $w$), and these singularities are converging on opposite sides of the contour $\gamma_{a}$ in the limit. We must manipulate the integrand accordingly.
@@
(106) $=\Phi_{\alpha,a}(w)^{-m^{\prime}}\phi_{a,N}(w)\tilde{\phi}_{+,a,N}(w)^{-1}$
@@
(111) $\tilde{\phi}_{-}(w)=\frac{1}{(1-w^{-1})^{N}}\Phi_{\alpha}(w)^{\frac{N}{2}}$
by the Weiner-Hopf factorization. [^14] See [3, Section 5] for details on how to compute this. Thus the correlation kernel for the split two-periodic Aztec diamond may be written in the following manner in the case where $0<m^{\prime}\leq\nicefrac{{N}}{{2}}$,
@@ 5.2.2. Proof when $m^{\prime}>\nicefrac{{N}}{{2}}$
We state the results of [3, Lemma 3.1] under the stated condition,
@@ 6. Analysis of the Eigen-Decomposition
We next need to inspect the possible orders of $x$ and $y$ that appear in the expansion. Notice that the exponent in the last sum is at most $j-2k+1$ and at least $-j+2k-1$. Moreover, the smallest exponent in the penultimate sum is $j$. Checking the possible powers on $x$ and $y$ gives,
@@ 6.2. Poles and Zeros of $r_{a,k,N}(w)$
(135) tr $\text{tr }\phi_{a,N}(w)=\left(\frac{1}{2}+g_{a,\alpha,\beta}(w)\right)\left(r_{a,\alpha,1}(w)^{\frac{N}{2}}r_{a,\beta,1}(w)^{\frac{N}{2}}+r_{a,\alpha,2}(w)^{\frac{N}{2}}r_{a,\beta,2}(w)^{\frac{N}{2}}\right)\\ +\left(\frac{1}{2}-g_{a,\alpha,\beta}(w)\right)\left(r_{a,\alpha,1}(w)^{\frac{N}{2}}r_{a,\beta,2}(w)^{\frac{N}{2}}+r_{a,\alpha,2}(w)^{\frac{N}{2}}r_{a,\beta,1}(w)^{\frac{N}{2}}\right)$
@@
any additional poles or zeros must also be poles of tr $\text{tr }\phi_{a,N}(w)$. Other than $w=a^{2}$, the potential poles of tr $\text{tr }\phi_{a,N}(w)$ come from the poles of $g_{a,\alpha,\beta}(w)$. These are explicitly,
@@
where coefficients $b_{a,-}$ and $b_{a,+}$ are also given in Appendix A. This gives us the information necessary to write the leading order terms of the Laurent expansion of tr $\text{tr }\phi_{a,N}(w)$ about $w=a^{2}$ and $w=a^{-2}$
(147) tr $\text{tr }\phi_{a,N}(w)=\frac{(c_{a,\alpha,+}c_{a,\beta,+})^{\frac{N}{2}}}{2(w-a^{2})^{2N}}+O\big{(}(w-a^{2})^{-2N+1}\big{)}$
(148) tr $\text{tr }\phi_{a,N}(w)=\frac{1}{2}(c_{a,\alpha,-}c_{a,\beta,-})^{\frac{N}{2}}+O\big{(}w-a^{-2}\big{)}$

### v5

@@ 4.4. Local Asymptotics of $I^{\alpha}_{2,1}$
From Lemma 4.2, we know that $\psi_{1}(z)$ has a maximum at $z_{1}^{*}$, so $\gamma_{0,1}$ should be deformed to the steepest ascent contour through this point, $\gamma_{s,1}$. After this deformation, we only need to show that
@@ 5. An Intermediate Correlation Kernel via a Non-Intersecting Paths Process
In Section 1.3, we introduced a non-intersection paths model that we claimed was equivalent to the split two-periodic Aztec diamond. In this section we begin by justifying the equivalence of the models. This explanation can be found in Section 5.1. These results are not new, the bijection was first detailed in [BD19, Section 5] and then further described in [CD23]. Once we introduce this process, the formulation of the kernel follows from a slight modification of [BD19, Theorem 5.2]. This work is done in Section 5.2.
@@ 5.1. Non-Intersecting Paths Process
Recall the definition of the Aztec diamond graph given by equations (1)-(3). From this definition, we would like to label the four types of edges found in the Aztec diamond graph. We call an edge of the form $((2j+1,2k),(2j+2,2k+1))$ an east edge, an edge of the form $((2j,2k+1),(2j,2k+2))$ a west edge, an edge of the form $((2j,2k+1),(2j+1,2k))$ a south edge, and an edge of the form ((2j+1,2k),(2j+2,2k1))((2j+1,2k),(2j+2,2k-1)) a north edge. The different types of edges are depicted in Figure 17.
@@ 5.1.1. The Aztec Diamond to DR Paths
We will start by relating the Aztec diamond graph to the DR lattice paths model, which was first introduced in [Jo03]. Under an appropriate choice of weighting this model is equivalent to the Aztec diamond. We will explain this bijection here. To start, we will describe the DR graph of size $n$. The DR graph of size $n$ has vertices given by,
@@
Figure 17. On the left is a covering of the Aztec diamond of size $n=4$. East, west, south, and north edges are depicted in blue, red, green, and yellow respectively. The corresponding paths on the DR graph are shown in the figure on the right.
@@ 5.1.2. The Tower Aztec Diamond Graph
The tower Aztec diamond graph was first described in [BD19] and later named in [CD23]. Informally, the $(n,p)$ tower Aztec diamond graph consists of two Aztec diamonds, one of size $n$ and the other of size n1n-1, stitched together by a strip of the rotated square grid of length $p$. An example of the tower Aztec diamond graph is show in Figure 18. The $(n,p)$ tower Aztec diamond graph has the vertex sets,
@@
[…]
One should realize that it is not possible to have a dimer configuration on the tower Aztec diamond where one dimer contains both a vertex from the corridor and a vertex from either of the appending Aztec diamonds. In fact, there is only one possible dimer configuration on the corridor. Thus the statistics of the tower Aztec diamond can be simplified to the statistics on the two appending Aztec diamond. See [CD23] for more detail.
@@ 5.1.3. Tower Aztec Diamond to Paths Process
The tower DR path process is nearly the process described in [BD19]. To get exactly the process in their work, we should change the diagonal edges of the DR tower graph into a combination of a horizontal edge and diagonal edge, as depicted in Figure 19. The additional horizontal edges will have weight one and thus not effect the statistics of the model. Once again we consider $n+p$ non intersection paths which start on the left at the top $n+p$ vertices and end on the right at the bottom $n+p$ vertices.
Thus if we send the length of the corridor to infinity, the correlation kernel for the top part of the model,referred to as $K_{top}$, in [BD19]) is also a correlation kernel of the Aztec diamond of size $n$. This correlation kernel is stated explicitly in [BD19, Theorem 3.1].
@@
Introducing this parameter $a\in(0,1)$ is necessary in order for us to directly apply [BD19, Theorem 3.1]. This is because the theorem does not permit singularities on the unit circle. We will call this weighting a split biased two-periodic weighting. Ultimately, we will send $a\to 1$ to obtain results for our model of interest. To compute the kernel, we will need to utilize the Weiner-Hopf factorization,
@@
For a full description of the derivation and properties of the matrices $\tilde{\phi}_{-,a,N}(z)$ and $\tilde{\phi}_{+,a,N}(z)$ see [BD19, Section 3]. It is important to note that $\tilde{\phi}_{+,a,N}(z)^{\pm 1}$ is analytic for $|z|<1$ and continuous for $|z|\leq 1$, while $\tilde{\phi}_{-,a,N}(z)^{\pm 1}$ is analytic for $|z|>1$ and continuous for $|z|\geq 1$. Moreover, the factorization commutes with the limit $a\to 1$.
@@ 5.2.1. Proof when $m^{\prime}\leq\nicefrac{{N}}{{2}}$
From [BD19, Theorem 3.1] we may immediately write the correlation kernel for the split biased two-periodic Aztec diamond,
@@
(102) $\phi_{m,a,N}(z)=\begin{cases}\Phi_{\alpha,a}(z)^{m}&\text{if }m\leq\nicefrac{{N}}{{2}}\\ \Phi_{\alpha,a}(z)^{\frac{N}{2}}\Phi_{\beta,a}(z)^{m-\frac{N}{2}}&\text{if }m>\nicefrac{{N}}{{2}}\end{cases}.$
The contour $\gamma_{a}$ is a simple closed curve around $a^{2}$, while the contour $\gamma_{0,1,a}$ is a simple closed curve containing 0, 11, and $\gamma_{a}$. To get the correlation kernel for the split two-periodic Aztec diamond, we wish to take the limit as $a\to 1$. The problem is the double contour integral in equation (101) is singular at both $a^{2}$ and $a^{-2}$ (with respect to $w$), and these singularities are converging on opposite sides of the contour $\gamma_{a}$ in the limit. We must manipulate the integrand accordingly. We have
@@
(104) $=\Phi_{\alpha,a}(w)^{-m^{\prime}}\phi_{a,N}(w)\tilde{\phi}_{+,a,N}(w)^{-1}.$
@@
(109) $\tilde{\phi}_{-,N}(w)=\frac{1}{(1-w^{-1})^{N}}\Phi_{\alpha}(w)^{\frac{N}{2}}$
by the Weiner-Hopf factorization. [^14] See [BD19, Section 5] for details on how to compute this. Thus the correlation kernel for the split two-periodic Aztec diamond may be written in the following manner in the case where $0<m^{\prime}\leq\nicefrac{{N}}{{2}}$,
@@ 5.2.2. Proof when $m^{\prime}>\nicefrac{{N}}{{2}}$
We state the results of [BD19, Lemma 3.1] under the stated condition,
@@ 6. Analysis of the Eigen-Decomposition
We next need to inspect the possible orders of $x$ and $y$ that appear in the expansion. Notice that the exponent in the last sum is at most j2k+1j-2k+1 and at least j+2k1-j+2k-1. Moreover, the smallest exponent in the penultimate sum is $j$. Checking the possible powers on $x$ and $y$ gives,
@@ 6.2. Poles and Zeros of $r_{a,k,N}(w)$
(133) $\text{tr }\phi_{a,N}(w)=\left(\frac{1}{2}+g_{a,\alpha,\beta}(w)\right)\left(r_{a,\alpha,1}(w)^{\frac{N}{2}}r_{a,\beta,1}(w)^{\frac{N}{2}}+r_{a,\alpha,2}(w)^{\frac{N}{2}}r_{a,\beta,2}(w)^{\frac{N}{2}}\right)\\ +\left(\frac{1}{2}-g_{a,\alpha,\beta}(w)\right)\left(r_{a,\alpha,1}(w)^{\frac{N}{2}}r_{a,\beta,2}(w)^{\frac{N}{2}}+r_{a,\alpha,2}(w)^{\frac{N}{2}}r_{a,\beta,1}(w)^{\frac{N}{2}}\right)$
@@
any additional poles or zeros must also be poles of $\text{tr }\phi_{a,N}(w)$. Other than $w=a^{2}$, the potential poles of $\text{tr }\phi_{a,N}(w)$ come from the poles of $g_{a,\alpha,\beta}(w)$. These are explicitly,
@@
where coefficients $b_{a,-}$ and $b_{a,+}$ are also given in Appendix A. This gives us the information necessary to write the leading order terms of the Laurent expansion of $\text{tr }\phi_{a,N}(w)$ about $w=a^{2}$ and $w=a^{-2}$
(145) $\text{tr }\phi_{a,N}(w)=\frac{(c_{a,\alpha,+}c_{a,\beta,+})^{\frac{N}{2}}}{2(w-a^{2})^{2N}}+O\big((w-a^{2})^{-2N+1}\big)$
(146) $\text{tr }\phi_{a,N}(w)=\frac{1}{2}(c_{a,\alpha,-}c_{a,\beta,-})^{\frac{N}{2}}+O\big(w-a^{-2}\big)$